# Query Performance Prediction using Relevance Judgments Generated by Large Language Models

### v2

@@ 7.2. Impact of fine-tuning and the choice of LLMs
To answer RQ4, we analyze the impact of fine-tuning and the choice of LLMs on the quality of generated relevance judgments and QPP. We consider LLaMA-7B (Touvron et al., 2023a), Llama-3-8B (AI@Meta, 2024), and Llama-3-8B-Instruct (AI@Meta, 2024) under two settings: (i) trained with PEFT on human relevance labels (following the same fine-tuning setup as in Section 5.6), and (ii) few-shot prompted (in-context learning) . [^6] We randomly sample human-labeled demonstration examples from the same set used for fine-tuning LLMs; each example is a triplet (query, passage, relevant/irrelevant); our experiments show that two examples work best, one with relevant passages and one with irrelevant passages. We do not report the results for a zero-shot setting because our preliminary experiments show that prompting these LLMs in a zero-shot way yields pretty poor performance.
@@
To evaluate the performance of judging relevance, we compute Cohen’s $\kappa$ metric to measure the agreement between relevance judgments made by the TREC assessors (i.e., relevance judgments in the qrels) and relevance judgments automatically generated by a fine-tuned or few-shot LLM, on TREC-DL 19–22. Faggioli et al. (2023a) reported the relevance judgment agreement in terms of Cohen’s $\kappa$ between TREC assessors and GPT-3.5 (text-davinci-003) on TREC-DL 21; we also consider their Cohen’s $\kappa$ value for comparison. To evaluate QPP quality, we compute the Pearson’s $\rho$ correlation coefficients between BM25’s actual nDCG@10 values and those predicted by QPP-GenRE using relevance judgments generated by an LLM, on TREC-DL 19–22.777We do not report the Pearson’s $\rho$ correlation for GPT-3.5 (text-davinci-003) because the relevance judgments generated by Faggioli et al. (2023a) are not available to us. The judging depth is set to 1000 in a ranked list. We report the results in Table 5.
@@
We have two observations. First, fine-tuning improves the quality of relevance judgment generation and QPP for all LLMs on all datasets. Specifically, all fine-tuned LLMs exhibit improved relevance judgment agreement with the TREC assessors on TREC-DL 19–22. After fine-tuning, LLaMA-7B and Llama-3-8B achieve “fair” agreement with the TREC assessors on TREC-DL 19, 20 and 21, [^8] Note that unlike the qrels files for TREC-DL 19, 20, and 21 which are fully manually annotated, the qrels file for TREC-DL 22 is constructed by first detecting near-duplicate items and manually judging only one representative item from each near-duplicate cluster for a given query (Craswell et al., 2021); this difference may result in variation in Cohen’s $\kappa$ values of LLMs across TREC-DL 19, 20, 21, and TREC-DL 22. Llama-3-8B-Instruct (fine-tuned) even achieves “moderate” agreement on TREC-DL 21 (a Cohen’s $\kappa$ value of 0.418). All fine-tuned LLMs exhibit a higher Cohen’s $\kappa$ value than the commercial LLM, GPT-3.5 (text-davinci-003). All fine-tuned LLMs surpass their corresponding few-shot counterpart on all datasets in terms of Pearson’s $\rho$. This reveals that fine-tuning is an effective way to improve the quality of LLMs in generating relevance judgments, which finally translates to better QPP quality.
@@
Second, newly released and instruction-tuned LLMs generally perform better. Llama-3-8B shows improved relevance judgment agreement and QPP quality compared to LLaMA-7B in most cases. Llama-3-8B-Instruct further enhances relevance judgment generation and QPP quality over both Llama-3-8B and LLaMA-7B across most cases. Notably, Llama-3-8B-Instruct (few-shot) even performs better than or equally as well as LLaMA-7B (fine-tuned) on TREC-DL 19 and 22. This finding implies that with a more effective LLM QPP-GenRE has the potential to achieve imrpoved QPP performance.
@@ 7.3. Integrating QPP-GenRE with an LLM-based re-ranker
To show QPP-GenRE’s compatibility with other types of relevance prediction methods instead of directly asking an LLM to explicitly generate explicit relevance judgments, we adapt a state-of-the-art pointwise LLM-based re-ranker, RankLLaMA (Ma et al., 2023a), into a relevance judgment generator, and then integrate QPP-GenRE with the adapted RankLLaMA. Specifically, we translate a re-ranking score into a relevance judgment by applying a threshold: an item is deemed as “relevant” if its re-ranking score meets or exceeds a given threshold value. We analyze Pearson’s $\rho$ and Kendall’s $\tau$ correlation coefficients between BM25’s actual nDCG@10 values and those predicted by QPP-GenRE integrated with RankLLaMA w.r.t. different threshold values on TREC-DL 19 and 20. We employ RepLLaMA (7B) from Tevatron. [^9] https://github.com/texttron/tevatron/tree/main/examples/rankllama RankLLaMA’s re-ranking scores for BM25 range from -12.93 to 89.90 for TREC-DL 19 and from -14.38 to 8.82 for TREC-DL 20. Thresholds are set at intervals of 0.5. The judging depth is set to 1,000 in a ranked list.
We report the results in Figure 4. We find that RepLLaMA achieves the highest QPP quality on both datasets when the threshold is 1. At this particular threshold, RepLLaMA achieves high Pearson’s $\rho$ values of 0.789 and 0.788 on TREC-DL 19 and 20, respectively. These values exceed those of fine-tuned LLaMA-7B, which achieves Pearson’s $\rho$ values of 0.715 and 0.627 on TREC-DL 19 and 20, respectively, as well as Llama-3-8B-Instruct, which achieves Pearson’s $\rho$ values of 0.647 and 0.743 on TREC-DL 19 and 20, respectively (see Figure 5). [^10] Note that the comparison is not fair because (i) LLaMA-7B, Llama-3-8B-Instruct and all other supervised QPP methods used in this paper are trained on the development set of MS MARCO V1, while RankLLaMA (Ma et al., 2023a) was trained on the training set of MS MARCO V1, which is much larger. (ii) We employ the official version of MS MARCO V1, while RankLLaMA (Ma et al., 2023a) uses the Tevatron version of MS MARCO V1, where passages are enriched with document titles; Lassance and Clinchant (2023) reveal that incorporating titles leads to enhanced ranking performance. This means that a state-of-the-art pointwise LLM-based re-ranker can be adapted into an effective relevance judgment generator. The high QPP quality achieved by RankLLaMA demonstrates QPP-GenRE’s compatibility with other types of relevance prediction methods besides directly using LLMs as relevance judgment generators (i.e., asking an LLM to explicitly generate explicit relevance judgments).
@@ 7.4. QPP-GenRE’s interpretability
As QPP-GenRE computes QPP based on generated relevance judgments, we analyze QPP errors from the perspective of relevance judgment generation. Figure 5 shows the QPP errors of QPP-GenRE integrated with LLaMA-7B in predicting the performance of BM25 and ANCE in terms of RR@10 on TREC-DL 19 and 20; the error is defined as the distance between the RR@10 values predicted by QPP-GenRE and actual RR@10 values, namely “predicted RR@10 minus actual RR@10.” We find that most RR@10 values predicted by QPP-GenRE tend to be smaller than the actual RR@10 values, indicating that QPP-GenRE performs less effectively in identifying relevant items than irrelevant ones in the top of the ranked list. Table 6 shows the confusion matrices that compare relevance judgments made by TREC assessors (i.e., relevance judgments in qrels) and QPP-GenRE integrated with LLaMA-7B on TREC-DL 19 and 20. We find that QPP-GenRE tends to wrongly predict some relevant items as irrelevant (false negatives), which provides a further interpretation of the QPP errors we found above. Therefore, reducing false negatives in generating relevance judgments is a potential way to improve the QPP quality of QPP-GenRE. We leave this exploration for future work.

### v3

@@ 7.2. Impact of fine-tuning and the choice of LLMs
To answer RQ4, we analyze the impact of fine-tuning and the choice of LLMs on the quality of generated relevance judgments and QPP. We evaluate two widely-used families of LLMs, Llama and Mistral, spanning from 1B to 70B under two settings: (i) trained with PEFT on human relevance labels (following the same fine-tuning setup as in Section 5.6), and (ii) few-shot prompted (in-context learning) . [^5] We randomly sample human-labeled demonstration examples from the same set used for fine-tuning LLMs; each example is a triplet (query, passage, relevant/irrelevant); our experiments show that two examples work best, one with relevant passages and one with irrelevant passages. For the Llama family, besides LLaMA-7B (Touvron et al., 2023), our evaluation includes Llama-3.2-1B-Instruct, Llama-3.2-3B-Instruct, Llama-3-8B, Llama-3-8B-Instruct, and Llama-3-70B-Instruct. For the Mistral family, we focus on Mistral-7B-Instruct-v0.3 and Mistral-22B-Instruct (a.k.a. Mistral-Small-Instruct-2409).
We do not report the results for a zero-shot setting because our preliminary experiments show that prompting these LLMs in a zero-shot way yields pretty poor performance. Note that we do not fine-tune Llama-3-70B-Instruct due to budget constraints.
@@
To evaluate the performance of judging relevance, we compute Cohen’s $\kappa$ metric to measure the agreement between relevance judgments made by the TREC assessors (i.e., relevance judgments in the qrels) and relevance judgments automatically generated by a fine-tuned or few-shot LLM, on TREC-DL 19–22. Faggioli et al. (2023a) reported the relevance judgment agreement in terms of Cohen’s $\kappa$ between TREC assessors and GPT-3.5 (text-davinci-003) on TREC-DL 21; we also consider their Cohen’s $\kappa$ value for comparison. To evaluate QPP quality, we compute the Pearson’s $\rho$ correlation coefficients between BM25’s actual nDCG@10 values and those predicted by QPP-GenRE using relevance judgments generated by an LLM, on TREC-DL 19–22.666We do not report the Pearson’s $\rho$ correlation for GPT-3.5 (text-davinci-003) because the relevance judgments generated by Faggioli et al. (2023a) are not available to us. The judging depth is set to 1000 in a ranked list. We show the results in Table 5 and Figure 5.
@@
We have three observations. First, fine-tuning generally markedly improves the quality of relevance judgment generation and QPP, particularly for LLMs with sizes ranging from 1 billion to 8 billion parameters. Specifically, almost all of fine-tuned LLMs exhibit improved relevance judgment agreement with the TREC assessors on TREC-DL 19–22. After fine-tuning, LLaMA-7B and Llama-3-8B achieve “fair” agreement with the TREC assessors on TREC-DL 19, 20 and 21, [^7] Note that unlike the qrels files for TREC-DL 19, 20, and 21 which are fully manually annotated, the qrels file for TREC-DL 22 is constructed by first detecting near-duplicate items and manually judging only one representative item from each near-duplicate cluster for a given query (Craswell et al., 2021); this difference may result in variation in Cohen’s $\kappa$ values of LLMs across TREC-DL 19, 20, 21, and TREC-DL 22. Llama-3-8B-Instruct (fine-tuned) even achieves “moderate” agreement on TREC-DL 21 (a Cohen’s $\kappa$ value of 0.418). All fine-tuned LLMs (except for Llama-3.2-1B-Instruct) exhibit a higher Cohen’s $\kappa$ value than the commercial LLM, GPT-3.5 (text-davinci-003). All fine-tuned LLMs (except for Mistral-22B-Instruct on TREC-DL 22) surpass their corresponding few-shot counterpart on all datasets in terms of Pearson’s $\rho$. This reveals that fine-tuning is an effective way to improve the quality of LLMs in generating relevance judgments, which finally translates to better QPP quality.
Second, larger LLMs with over 22 billion parameters demonstrate significantly greater effectiveness than their smaller counterparts in the few-shot setting. Specifically, in this setting, Llama-3-70B-Instruct achieves the best overall performance. Mistral-22B-Instruct consistently outperforms Mistral-7B-Instruct-v0.3 across all datasets. However, a fine-tuned Llama model with only 3 billion parameters markedly outperforms both of these larger few-shot LLMs across all datasets.
@@
Third, Instruction-tuned LLMs generally perform better than their standard counterparts. Llama-3-8B-Instruct further enhances relevance judgment generation and QPP quality over both Llama-3-8B and LLaMA-7B across most cases. Notably, Llama-3-8B-Instruct (few-shot) even performs better than or equally as well as LLaMA-7B (fine-tuned) on TREC-DL 19 and 22. This finding implies that with a more effective LLM QPP-GenRE has the potential to achieve improved QPP performance.
The above observations provide insights into the minimum requirements needed to achieve reliable QPP quality. Our findings show that a fine-tuned 3B model (Llama-3.2-3B-Instruct) offers the best trade-off between QPP quality and computational overhead: it not only markedly outperforms few-shot 70B LLMs, but also delivers QPP quality comparable to that of fine-tuned 7/8B LLMs.
@@ 7.3. Integrating QPP-GenRE with an LLM-based re-ranker
To show QPP-GenRE’s compatibility with other types of relevance prediction methods instead of directly asking an LLM to explicitly generate explicit relevance judgments, we adapt a state-of-the-art pointwise LLM-based re-ranker, RankLLaMA (Ma et al., 2023a), into a relevance judgment generator, and then integrate QPP-GenRE with the adapted RankLLaMA. Specifically, we translate a re-ranking score into a relevance judgment by applying a threshold: an item is deemed as “relevant” if its re-ranking score meets or exceeds a given threshold value. We analyze Pearson’s $\rho$ and Kendall’s $\tau$ correlation coefficients between BM25’s actual nDCG@10 values and those predicted by QPP-GenRE integrated with RankLLaMA w.r.t. different threshold values on TREC-DL 19 and 20. We employ RankLLaMA (7B) from Tevatron. [^8] https://github.com/texttron/tevatron/tree/main/examples/rankllama RankLLaMA’s re-ranking scores for BM25 range from -12.93 to 89.90 for TREC-DL 19 and from -14.38 to 8.82 for TREC-DL 20. Thresholds are set at intervals of 0.5. The judging depth is set to 1,000 in a ranked list.
We report the results in Figure 6. We find that RankLLaMA achieves the highest QPP quality on both datasets when the threshold is 1. At this particular threshold, RankLLaMA achieves high Pearson’s $\rho$ values of 0.789 and 0.788 on TREC-DL 19 and 20, respectively. These values exceed those of fine-tuned LLaMA-7B, which achieves Pearson’s $\rho$ values of 0.715 and 0.627 on TREC-DL 19 and 20, respectively, as well as Llama-3-8B-Instruct, which achieves Pearson’s $\rho$ values of 0.647 and 0.743 on TREC-DL 19 and 20, respectively (see Figure 5). [^9] Note that the comparison is not fair because (i) LLMsand all other supervised QPP methods used in this paper are trained on the development set of MS MARCO V1, while RankLLaMA (Ma et al., 2023a) was trained on the training set of MS MARCO V1, which is much larger. (ii) We employ the official version of MS MARCO V1, while RankLLaMA (Ma et al., 2023a) uses the Tevatron version of MS MARCO V1, where passages are enriched with document titles; Lassance and Clinchant (2023) reveal that incorporating titles leads to enhanced ranking performance. This means that a state-of-the-art pointwise LLM-based re-ranker can be adapted into an effective relevance judgment generator. The high QPP quality achieved by RankLLaMA demonstrates QPP-GenRE’s compatibility with other types of relevance prediction methods besides directly using LLMs as relevance judgment generators (i.e., asking an LLM to explicitly generate explicit relevance judgments).
@@
Table 6. Correlation coefficients (Pearson’s $\rho$ and Kendall’s $\tau$) between actual retrieval quality, in terms of nDCG@3, of ConvDR (Yu et al., 2021) and its performance predicted by QPP-GenRE/baselines, on CAsT-19 and 20. Following Meng et al. (2023c), for QPP methods requiring queries as input, we feed them with either T5-generated or human-written query rewrites. indicates statistically significant correlation coefficients ($p$-value $<0.05$). indicates the statistically significant improvement of QPP-GenRE ($n$=200) compared to all the baselines (paired $t$-test; $p$-value $<0.001$ with Bonferroni correction for multiple testing). The best value in each column is marked in bold. $n$ denotes QPP-GenRE’s judgment depth in a ranked list.
Ranker: ConvDR
T5-generated query rewrites Human-written query rewrites
CAsT-19 CAsT-20 CAsT-19 CAsT-20
QPP method P-$\rho$ K-$\tau$ P-$\rho$ K-$\tau$ P-$\rho$ K-$\tau$ P-$\rho$ K-$\tau$
Clarity 0.257 0.176 0.126 0.088 0.257 0.176 0.126 0.088
WIG 0.387 0.274 0.377 0.277 0.412 0.285 0.384 0.264
NQC 0.431 0.307 0.339 0.261 0.431 0.307 0.339 0.261
$\sigma_{max}$ 0.378 0.267 0.282 0.219 0.378 0.267 0.282 0.219
n($\sigma_{x\%}$) 0.187 0.175 0.199 0.168 0.216 0.196 0.201 0.156
SMV 0.386 0.285 0.275 0.216 0.386 0.285 0.275 0.216
UEF(NQC) 0.435 0.312 0.343 0.265 0.427 0.310 0.341 0.263
RLS(NQC) 0.429 0.311 0.337 0.267 0.413 0.308 0.342 0.259
QPP-PRP 0.350 0.270 0.280 0.210 0.345 0.265 0.275 0.205
NQA-QPP 0.175 0.115 0.082 0.075 0.142 0.091 0.065 0.058
BERTQPP 0.243 0.170 0.236 0.185 0.256 0.172 0.262 0.209
qppBERT-PL 0.203 0.169 0.181 0.165 0.105 0.090 0.166 0.161
M-QPPF 0.242 0.174 0.285 0.219 0.262 0.190 0.313 0.254
QPP-GenRE ($n=200$) 0.623 0.505 0.484 0.395 0.645 0.529 0.678 0.551
QPP-GenRE ($n=10$) 0.617 0.504 0.471 0.388 0.619 0.506 0.659 0.534
QPP-GenRE ($n=100$) 0.623 0.505 0.485 0.396 0.644 0.529 0.675 0.547
QPP-GenRE ($n=1,000$) 0.623 0.505 0.487 0.398 0.645 0.529 0.684 0.556
7.4. Generalization to conversational search
To assess the generalizability of QPP-GenRE to new domains, we apply it to the conversational search scenario (Mo et al., 2023a, b, 2024b, 2024a, 2025; Meng et al., 2025b) in a zero-shot manner. Specifically, we evaluate QPP-GenRE and other baselines on predicting the performance of ConvDR (Yu et al., 2021), a widely used conversational dense retriever. Given the findings in Section 7.2, which show that the fine-tuned Llama-3.2-3B-Instruct model achieves high relevance prediction quality at low inference cost, we equip QPP-GenRE with this model fine-tuned on MS MARCO V1 for relevance prediction. For all supervised QPP baselines, we directly use their checkpoints trained on MS MARCO V1 for assessing ANCE (see Section 6.2). Because a user query in a conversation depends on the conversational context, i.e., a query may contain omissions, coreferences, or ambiguities, it is hard for existing QPP methods to capture users’ information need from such context-dependent queries. Therefore, we follow Meng et al. (2023c) to provide QPP methods (including QPP-GenRE) with self-contained query rewrites as input. These rewrites are either generated by the T5 query generator [^10] https://huggingface.co/castorini/t5-base-canard or written by humans. Table 6 presents the performance of QPP-GenRE along with all the baselines on assessing ConvDR (Yu et al., 2021) in terms of nDCG@3 on the CAsT-19 (Dalton et al., 2020b) and 20 (Dalton et al., 2020a) datasets. Note that nDCG@3 is the primary evaluation metric officially adopted by TREC CAsT (Dalton et al., 2020b). We have two main observations.
First, QPP-GenRE significantly outperforms all QPP baselines on both datasets when provided with either type of query input. This demonstrates the ability of QPP-GenRE to generalize effectively to the conversational search domain. Second, QPP-GenRE achieves higher performance when provided with human-written query rewrites compared to T5-generated rewrites on both datasets. This finding highlights the critical role of high-quality query rewrites in effectively adapting QPP methods to conversational search scenarios; this finding also aligns with prior research (Meng et al., 2023c).
@@ 7.5. QPP-GenRE’s interpretability
Table 8. Performance of each class for QPP-GenRE with LLaMA-7B on the TREC-DL 2019 and 2020 qrels.
TREC-DL 19 TREC-DL 20
Class Precision Recall F1 Precision Recall F1
Relevant 0.576 0.301 0.395 0.389 0.292 0.333
Irrelevant 0.780 0.918 0.844 0.884 0.922 0.902
Table 9. Retrieval quality, in terms of RR@10, predicted by various QPP methods for the BM25 retrieval of the query “who is Robert Gra” (query ID 1037798) on TREC-DL 19.
QPP Methods Retrieval quality
WIG 2.427
NQC 0.106
BERTQPP 0.172
qppBERT-PL 0.368
QPP-GenRE 0.500
Ground-truth RR@10 1.000
Table 10. Ranked list returned by BM25 for the query “who is Robert Gra” (query ID 1037798), human-labeled relevance judgments and ones predicted by QPP-GenRE with LLaMA-7B on TREC-DL 19.
Rank Passage Human QPP-GenRE
1 Captain Robert Gray, May 1972. Discovering the Columbia River, May 1792 … The Columbia River was given the name it bears today in May 1792… Relevant Irrelevant
2 Robert Gray. A surprise came on the Democratic side in the race for Mississippi Governor. Robert Gray, a retired firefighter and truck driver… Irrelevant Relevant
3 Team Mississippi Robert Gray For Governor Official Page. Robert Gray never would have made it without God… Irrelevant Irrelevant
4 I’m not a politician, said Gray in a Wednesday interview. I’m not a person who really wanted to run for Governor. Robert Gray is a 46-year-old truck driver… Irrelevant Relevant
As QPP-GenRE computes QPP based on generated relevance judgments, we analyze QPP errors from the perspective of relevance judgment generation. Figure 7 shows the QPP errors of QPP-GenRE integrated with LLaMA-7B in predicting the performance of BM25 and ANCE in terms of RR@10 on TREC-DL 19 and 20; the error is defined as the distance between the RR@10 values predicted by QPP-GenRE and actual RR@10 values, namely “predicted RR@10 minus actual RR@10.” We find that most RR@10 values predicted by QPP-GenRE tend to be smaller than the actual RR@10 values, indicating that QPP-GenRE performs less effectively in identifying relevant items than irrelevant ones in the top of the ranked list. Table 7 shows the confusion matrices that compare relevance judgments made by TREC assessors (i.e., relevance judgments in qrels) and QPP-GenRE integrated with LLaMA-7B on TREC-DL 19 and 20. Table 8 provides a detailed breakdown of QPP-GenRE’s prediction performance for each class, including metrics such as Precision, Recall, and F1 score. We find that QPP-GenRE tends to wrongly predict some relevant items as irrelevant (false negatives), which provides a further interpretation of the QPP errors we found above. Therefore, reducing false negatives in generating relevance judgments is a potential way to improve the QPP quality of QPP-GenRE. We leave this exploration for future work.
To show the superior interpretability of QPP-GenRE compared to other baselines, we provide a case study shown in Tables 9 and Tables 10. Table 9 lists the predicted or ground-truth retrieval quality in terms of RR@10 of BM25 for the query “who is Robert Gra” on TREC-DL 2019. The predictions are made using widely-used unsupervised QPP methods (WIG, NQC), supervised QPP methods (BERTQPP, qppBERT-PL), and QPP-GenRE.
We observe that the ground-truth RR@10 score for the query is 1, while the score predicted by QPP-GenRE is 0.5. QPP-GenRE infers RR@10 directly from the predicted relevance judgments for items in BM25’s ranked list, we can conclude that while the top-ranked item is actually relevant to the query, QPP-GenRE mistakenly classified it as irrelevant. Table 10 provides supporting evidence by displaying the relevance judgments assigned by human annotators and QPP-GenRE for each item in BM25’s ranked list. We found that QPP-GenRE fails to identify the top-ranked item as relevant. Specifically, QPP-GenRE does not identify the key part “Captain Robert Gray” in this item. This suggests that we could potentially improve QPP-GenRE’s performance by further fine-tuning it to predict relevance on query–item pairs specifically related to influential historical figures.
However, all baselines lack interpretability compared to QPP-GenRE. WIG and NQC do not directly predict values for a specific IR metric, and their scores are difficult to interpret in isolation without comparing them to the scores for other queries. BERTQPP is trained to predict RR@10, but in this case, it returns a score of 0.3, indicating inaccurate performance prediction. Unfortunately, BERTQPP does not provide any intermediate outputs to help understand why it made this error or how its performance can be improved, limiting its interpretability and actionable insights. qppBERT-PL first predicts the number of relevant items in each chunk of the ranked list and then aggregates these numbers into an overall score. While it is possible to check the intermediate predictions of the number of relevant items per chunk, this information is too coarse to provide detailed insights. In contrast, QPP-GenRE predicts the relevance of each individual item, offering more granular and informative insights.